2. Let X1, X2, X3 ..., X, be iid b(1, p) random variables. Let Sn = 27-1Xthen prove that Sn-E(Sn) N(0,1) as n +00. (Sn)
Answer #1

Golno we have xi v B C I J P) then Ecais =P Vcxij = P9 on: IX. BCDP) then Econ) = np. v (sn). - npgford consider, z = SD- ECOtan eropa - tnp e unpa ton OP9 -tne e unpa . E Σι; = BE orapa stomer . Mixi (t/1opq.) e tempat Mxi.Chunpg) - to Munct) = etapMai (Free) = = 1+ (women xi) + b (dowa 21) (Compt 2)... Mxi 1 + npa Eco). ) ECB) * ( Mxit ) = top 1+ t illit t?e2+ t Unpa anpwith in where ones than or Wheze, O. is higher order term power of in greater than or equal to 3/2 denominator log Mxi (Inpg)Tin log No CE) hingga 109 Mzo C+) = 2 100 log Man.ct) = /2t² s as nooo. ... Men (t) = eltz el Hence, . This by uniqueness pro